# GIFT: Unlocking Full Potential of Labels in Distilled Dataset at Near-zero Cost

### v1

@@ Abstract
Recent advancements in dataset distillation have demonstrated the significant benefits of employing soft labels generated by pre-trained teacher models. In this paper, we introduce a novel perspective by emphasizing the full utilization of labels. We first conduct a comprehensive comparison of various loss functions for soft label utilization in dataset distillation, revealing that the model trained on the synthetic dataset exhibits high sensitivity to the choice of loss function for soft label utilization. This finding highlights the necessity of a universal loss function for training models on synthetic datasets. Building on these insights, we introduce an extremely simple yet surprisingly effective plug-and-play approach, GIFT, which encompasses soft label refinement and a cosine similarity-based loss function to efficiently leverage full label information. Extensive experiments demonstrate that GIFT consistently enhances the state-of-the-art dataset distillation methods across various scales datasets without incurring additional computational costs. For instance, on ImageNet-1K with IPC = 10, GIFT improves the SOTA method RDED by 3.9% and 1.8% on ConvNet and ResNet-18, respectively. Code: https://github.com/LINs-lab/GIFT.  [^1] We have provided Pytorch implementation code in Appendix E
@@ 1 Introduction
Dataset distillation (DD) [33] has demonstrated its potential to significantly reduce data size while maintaining comparable model performance [3, 4, 42, 32, 41, 44]. Most existing DD methods focus on optimizing images [33, 42, 40, 3, 32, 14, 18, 41], but recent studies [36, 28, 25, 11] have highlighted the substantial benefits of soft labels. These studies utilize labels assigned by pre-trained models (also called teacher models), yielding siginificant enhancement in stability during the synthesizing process and considerable performance. The success can be attributed to their rich intra-class information, which is particularly valuable for training on extremely small datasets [28].
@@
To fully explore the utilization of soft labels in current state-of-the-art dataset distillation methods, we conduct a comprehensive comparison of loss functions for soft label utilization in synthesic datasets via the SOTA dataset distillation methods on Tiny-ImageNet and ImageNet-1K  [^2] The synthetic dataset of DATM for ImageNet-1k is not provided, so comparing with DATM is not possible on this dataseton $\texttt{IPC}\in\{1,10,50\}$  [^3] Detailed experiments are elaborated in Section 3. The results of Tiny-ImageNet and ImageNet-1K when IPC = 1 and IPC = 50 are shown in Appendix D. As illustrated in Figure 1, different dataset distillation methods employ distinct loss functions, and performance varies significantly across these loss functions. In particular, simply replacing KL divergence loss [13] with Soft cross-entropy loss [2] for SRe2L results in a significant performance drop of 13.3% on Tiny-ImageNet. This substantial gap indicates that models trained on synthetic datasets are highly sensitive to the choice of loss function. Hence, it is crucial to propose a universal and effective loss function that is robust across various synthetic datasets.
@@
Moreover, one challenge of soft labels is that they are inherently suboptimal, as the performance of the teacher model itself may be limited. For instance, a teacher model trained on ImageNet-1k on ConvNet has only 43.6% accuracy, a common model used in [28, 36] to assign labels. Furthermore, in practical scenarios, it is common for the teacher model to have a smaller architecture than the student model due to cross-architecture challenge [28]. This discrepancy further limits the performance of the student model when only relying on an inferior teacher model. By contrast, hard labels are accurate and provide reliable supervision information. However, directly utilizing hard labels through cross-entropy loss is not effective, as demonstrated in previous studies [5] and our experiments in Section 5.6. Therefore, it is imperative to effectively integrate hard label information to mitigate the limitations associated with soft labels.
@@
  1. 1.
    To the best of our knowledge, this paper is the first to provide a comprehensive comparison of loss functions for label utilization in dataset distillation.
  2. 2.
    Our study reveals the intriguing fact that models trained on synthetic datasets are highly sensitive to the choice of loss function.
  3. 3.
    We propose GIFT, a simple and universal label utilization algorithm for dataset distillation. GIFT is built on top of the off-the-shelf dataset distillation methods and requires no extra information, thus raising no additional cost.
  4. 4.
    Experiments verify that GIFT significantly improves performance over the state-of-the-art dataset distillation methods across varying scales and resolutions. It also enhances cross-architecture and cross-optimizer generalization.
(a) Tiny-ImageNet
(b) ImageNet-1K
Figure 1: Top-1 accuracy on various synthetic datasets via the SOTA dataset distillation methods across loss functions on Tiny-ImageNet and ImageNet-1K when IPC=10. value means the results of the loss function used by the distillation method itself (e.g., SRe2L [36] and RDED [28] use KL divergence loss [13]). value means the results of our proposed method GIFT, and ($\uparrow$) denotes improvements over the dataset distillation methods. It is obvious that our method GIFT significantly enhances the dataset distillation methods.
@@ 2 Related Work
In the field of dataset distillation, the majority of existing methods fix the labels as a one-hot format [33, 42, 40, 3, 32, 14, 18, 41], with a primary focus on optimizing synthetic images. Recent research highlights the significant benefits of utilizing soft labels to enhance model performance [28, 11, 36]. These benefits are particularly pronounced when training models on extremely small synthetic datasets, due to the enriched intra-class information soft labels provide. Methods for obtaining soft labels can be broadly classified into two categories.
@@
The first type involves learning labels, with several studies [1, 20, 27, 46, 7, 11] finding that making labels learnable can significantly improve performance and efficiency. Recent work [11] further highlights that optimizing labels can enhance training stability and improve performance.
@@
The subsequent trajectory of works seeks to directly obtain soft labels. Inspired by knowledge distillation [13], TESLA [5]introduces a soft label assignment strategy, directly generating soft labels by leveraging pre-trained teacher models trained on real datasets. These soft labels provide rich intra-class information, thereby improving distillation performance. Following this trend, the state-of-art methods [36, 28, 25] also utilize soft labels predicted by teacher models, achieving significant improvements.
@@ 3.1 Preliminary
Given a large dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the input sample and $y_{i}\in\{1,\ldots,C\}$ denotes the corresponding hard label, the objective of dataset distillation is to generate a synthetic dataset $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},\tilde{y}_{j})\}_{j=1}^{M}$, such that a model trained on $\mathcal{S}$ performs comparably to one trained on $\mathcal{D}$. In this paper, we explore how to fully utilize both soft labels and hard labels in given datasets. Therefore, we re-define the synthetic dataset $\mathcal{S}$ as $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},y_{j},\tilde{y}_{j})\}_{j=1}^{M}$, where $\tilde{\mathbf{x}}_{j}$ denotes the synthetic images, and $y_{j}$ and $\tilde{y}_{j}$ denote the corresponding hard labels and soft labels, respectively.
@@ 3.2 Are Loss Functions Pulling the Strings in Synthetic Dataset Performance?
Why do we need to answer this question? Labels in dataset distillation are commonly and typically utilized through a variety of established loss functions, such as cross-entropy (CE) [2], Kullback-Leibler (KL) divergence [13], soft cross-entropy [2] or mean squared error (MSE) [21]. Specifically, SRe2L [36] and RDED [28] employ KL divergence, DATM [11] uses soft cross-entropy, and G-VBSM [25] simultaneously utilizes MSE and CE. However, different distillation methods employ varying loss functions to train models on synthetic datasets, yet there is a notable lack of comprehensive comparison among these loss functions. Hence, we investigate the performance of four state-of-the-art dataset distillation methods under three commonly used loss functions.
Experiment settings. Our experiments span small-scale dataset Tiny-ImageNet and large-scale dataset ImageNet-1K on $\texttt{IPC}\in\{1,10,50\}$ using ConvNet [11]. Note that the synthetic dataset for DATM on ImageNet-1K is unavailable, precluding comparisons on this dataset. We evaluate at $\texttt{IPC}\in\{1,10,50\}$. Additional visualizations for Tiny-ImageNet and ImageNet-1K at $\texttt{IPC}=1$ and $IPC=50$ are provided in Appendix D. Notably, our empirical study is conducted from the end-user perspective: we treat the distillation of the synthetic dataset as a black box and apply different loss functions during the evaluation of the synthetic datasets.
Analysis. The results, visualized in Figure 1, reveal that even for the same synthetic dataset, the evaluation model performance varies significantly under different loss functions. For example, the performance of SRe2L decreases by 13.2% on Tiny-ImageNet when switching from KL loss to soft CE loss but improves by 0.7% when using MSE+CE loss. These findings illustrate that the performance of models trained on synthetic datasets is highly sensitive to the choice of the loss function, highlighting the necessity for a unified and effective loss function in dataset distillation.
@@ 4 Method
GIFT includes two key modules, namely, label refinement and a cosine similarity-based loss function. The former aims to refine soft labels by incorporating an additional smoothing label derived from hard label smoothing [29], while the latter is theoretically validated to achieve optimal performance simply using cosine similarity as the loss function. A PyTorch implementation of our method is provided in Appendix E.
@@
The refined soft label is defined as $\tilde{y}_{j}\leftarrow\gamma\cdot\frac{y_{i}}{\|y_{j}\|}+(1-\gamma)\cdot\frac{\tilde{y}_{j}}{\|\tilde{y}_{j}\|}$, where $j$ is the j-th synthetic images, $y_{j}$ is the smoothed label obtained via label smoothing technique for hard label, and $\tilde{y}_{j}$ is the soft label. In our experiments, $\gamma=0.1$ is validated to be optimal through experiments depicted in Figure 3 and Figure 8 in Appendix D.
@@
Prior study [28] has posited that representation learning from any samples $X$ to targets $Y$ is based on maximizing their mutual information. They propose to distill the dataset by maximizing $I_{\mathcal{V}}(X,Y)$, where $I_{\mathcal{V}}$ denotes the $\mathcal{V}$-information [34], which has demonstrated superior performance. However, we observe that most prior studies, including [28], have not investigated training models using $I_{\mathcal{V}}(X,Y)$. To address this gap, we explore the application of $I_{\mathcal{V}}(X,Y)$ by deriving an upper bound for the $\mathcal{V}$-information, as presented in Theorem 1 (see Appendix A for the proof).
Theorem 1
@@
where $\tau$ denotes the temperature parameter, $\mathcal{L}_{\textup{InfoNCE}}$ [22] serves as a proxy for $I_{\mathcal{V}}(X,Y)$ [28], and $f$ represents the similarity function.
Moreover, the targets $Y$ in the distilled dataset are pre-generated using pre-trained models [28, 36]. These targets can be considered high-dimensional vectors that are approximately orthogonal to each other. Consequently, the term ${\mathbb{E}}[\boldsymbol{\phi}_{\boldsymbol{\theta}}(\mathbf{x}_{i})\cdot y_{j}]$ approaches zero as $\mathbb{E}\left[\frac{\boldsymbol{\phi}_{\boldsymbol{\theta}}(\mathbf{x}_{i})\cdot y_{i}}{\|\boldsymbol{\phi}_{\boldsymbol{\theta}}(\mathbf{x}_{i})\|\|y_{i}\|}\right]$ approaches one, allowing (1) to be simplified as:
@@ 5 Experiments
In this section, we evaluate the superiority of our proposed GIFT across various datasets and neural architectures. First, we demonstrate the superior improvements of GIFT over the state-of-the-art dataset distillation in Section 5.2 and knowledge distillation methods in Section 5.3. Next, we verify that our proposed GIFT significantly enhances cross-architecture and cross-optimizer generalization in Section 5.4, and improves continual learning in Section 5.5. Furthermore, we conduct extensive ablation experiments to investigate the effect of each component of our method in Section 5.6. More details about experiments details and results can be found in Appendix C and Appendix D.

### v2

@@ Abstract
Recent advancements in dataset distillation have demonstrated the significant benefits of employing soft labels generated by pre-trained teacher models. In this paper, we introduce a novel perspective by emphasizing the full utilization of labels. We first conduct a comprehensive comparison of various loss functions for soft label utilization in dataset distillation, revealing that the model trained on the synthetic dataset exhibits high sensitivity to the choice of loss function for soft label utilization. This finding highlights the necessity of a universal loss function for training models on synthetic datasets. Building on these insights, we introduce an extremely simple yet surprisingly effective plug-and-play approach, GIFT, which encompasses soft label refinement and a cosine similarity-based loss function to efficiently leverage full label information. Extensive experiments indicate that GIFT consistently enhances state-of-the-art dataset distillation methods across various dataset scales, without incurring additional computational costs. Importantly, GIFT significantly enhances cross-optimizer generalization, an area previously overlooked. For instance, on ImageNet-1K with IPC = 10, GIFT enhances the state-of-the-art method RDED by 30.8% in cross-optimizer generalization  [^1] Our code is available at https://github.com/LINs-lab/GIFT. We also have provided Pytorch implementation code in Appendix F..
@@ 1 Introduction
Dataset distillation (DD) (Wang et al., 2018) has demonstrated its potential to significantly reduce data size while maintaining comparable model performance (Cazenavette et al., 2022; 2023; Zhao et al., 2020; Zhao & Bilen, 2023; Zhao et al., 2023). Most existing DD methods focus on optimizing images (Wang et al., 2018; Zhao & Bilen, 2022; Cazenavette et al., 2022; Kim et al., 2022; Liu et al., 2022), but recent studies (Yin et al., 2023; Sun et al., 2024; Shao et al., 2024; Guo et al., 2024) have highlighted the substantial benefits of soft labels. These studies utilize labels assigned by pre-trained models (also called teacher models), yielding siginificant enhancement in stability during the synthesizing process and considerable performance. Moreover, a notable study by Qin et al. (2024) examines the role of soft labels in depth, emphasizing their importance.
@@
To fully explore the utilization of soft labels in the state-of-the-art dataset distillation methods, we conduct a comprehensive comparison of loss functions for soft label in synthesic datasets of $\texttt{IPC}=10$  [^2] Detailed experiments are elaborated in Section 3. via the SOTA dataset distillation methods on Tiny-ImageNet and large-scale ImageNet-1K  [^3] The synthetic dataset of DATM for ImageNet-1k is not provided, so comparing with DATM is impossible.. As shown in Figure 1, different dataset distillation methods employ distinct loss functions, and performance varies significantly across loss functions. In particular, simply replacing KL divergence loss (Hinton et al., 2015) with Soft cross-entropy loss (Bridle, 1989) for SRe2L results in a significant performance drop of 13.3% on Tiny-ImageNet. This substantial gap indicates that models trained on synthetic datasets are highly sensitive to the choice of loss function.
Furthermore, we observe a notable performance degradation in these loss functions when applied across different optimizers. For example, when utilizing distilled data produced by RDED and training with the KL divergence loss, altering the optimizer from AdamW  [^4] RDED employs AdamW (Loshchilov, 2017) for evaluation. to Adam results in a performance decrease from 47.5% to 17.8% (as shown in Table 8). Hence, it is crucial to propose a universal and effective loss function that is robust across various scenarios.
(a) Tiny-ImageNet
(b) ImageNet-1K
Figure 1: Top-1 accuracy on various synthetic datasets via the SOTA dataset distillation methods across loss functions on Tiny-ImageNet and ImageNet-1K when IPC =10. value means the results of the loss function used by the distillation method itself (e.g., SRe2L (Yin et al., 2023) uses KL divergence (Hinton et al., 2015)). value means the results of our GIFT, and ($\uparrow$) denotes improvements over the dataset distillation methods. It is obvious that our method GIFT significantly enhances the dataset distillation methods.
@@
Moreover, one challenge of soft labels is that they are inherently suboptimal, as the performance of the teacher model itself may be limited. For instance, a teacher model trained on ImageNet-1k on ConvNet has only 43.6% test accuracy. Despite its worse performance, such model is frequently used in Sun et al. (2024); Yin et al. (2023) to assign labels. Furthermore, in practical scenarios, it is common for the teacher model to have a smaller architecture than the student model due to cross-architecture challenge (Sun et al., 2024), which further limits the performance of the student model when only relying on an inferior teacher model. By contrast, hard labels are accurate and provide reliable supervision information. Nonetheless, directly utilizing hard labels through cross-entropy loss is not effective, as demonstrated in previous studies (Cui et al., 2023) and our experiments in Section 5.6. Therefore, it is crucial to effectively integrate hard label information to mitigate the limitations associated with soft labels.
@@
  1. (a)
    To the best of our knowledge, this paper is the first to provide a comprehensive comparison of loss functions for label utilization in dataset distillation. Our study reveals the intriguing fact that models trained on synthetic datasets are highly sensitive to the choice of loss function.
  2. (b)
    We propose GIFT, a simple and universal label utilization algorithm including label refinement and a cosine similarity-based loss function. GIFT is built on top of the off-the-shelf dataset distillation methods and requires no extra information, thus raising no additional cost. Moreover, we provide a theoretical analysis to support the proposed use of cosine similarity.
  3. (c)
    We identify a critical issue that has been overlooked in prior research: cross-optimizer generalization, as defined in Section 3.1. We reveal that traditional loss functions suffer from significant robustness deficiencies when applied across different optimizers as detailed in Section 5.5. In contrast, GIFT significantly enhances dataset distillation methods in cross-optimizer generalization. We conduct both empirical and theoretical analyses of this challenge in Appendix E.
  4. (d)
    Experiments demonstrate that GIFT significantly improves performance over the state-of-the-art dataset distillation methods across varying scales and resolutions datasets, particularly for large-scale dataset distillation tasks. Furthermore, GIFT significantly enhances dataset distillation methods in cross-architecture, cross-optimizer generalization and proves advantageous in applications such as continual learning.
@@ 2 Related Work
In dataset distillation, the majority of existing methods fix the labels as a one-hot format (Wang et al., 2018; Zhao & Bilen, 2022; Cazenavette et al., 2022; Liu et al., 2022; Zhao & Bilen, 2023), with a primary focus on optimizing synthetic images. Recent research highlights the significant benefits of utilizing soft labels to enhance model performance (Sun et al., 2024; Guo et al., 2024; Yin et al., 2023). Methods for obtaining soft labels can be broadly classified into two categories.
@@
The first type involves learning labels, with several studies (Bohdal et al., 2020; Nguyen et al., 2020; Sucholutsky & Schonlau, 2021; Zhou et al., 2022; Guo et al., 2024) finding that learning labels can significantly improve performance. Recent work (Guo et al., 2024) highlights that optimizing labels can enhance training stability and improve performance.
@@
The subsequent works directly obtain soft labels. Inspired by knowledge distillation (Hinton et al., 2015), TESLA (Cui et al., 2023)introduces a soft label assignment strategy, directly generating soft labels by leveraging pre-trained teacher models trained on real datasets. These soft labels provide rich intra-class information, thereby improving distillation performance. Following this trend, the state-of-art methods (Yin et al., 2023; Sun et al., 2024; Shao et al., 2024) also utilize soft labels predicted by teacher models, achieving significant improvements.
@@ 3.1 Preliminary
Dataset Distillation.
Given a large dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the input sample and $y_{i}\in\{1,\ldots,C\}$ denotes hard label, the objective of dataset distillation is to generate a synthetic dataset $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},\tilde{y}_{j})\}_{j=1}^{M}$, such that a model trained on $\mathcal{S}$ performs comparably to one trained on $\mathcal{D}$. We explore how to fully utilize both soft labels and hard labels in given datasets. Therefore, we re-define the synthetic dataset $\mathcal{S}$ as $\mathcal{S}=\{(\tilde{\mathbf{x}}_{j},y_{j},\tilde{y}_{j})\}_{j=1}^{M}$, where $\tilde{\mathbf{x}}_{j}$ denotes the synthetic images, and $y_{j}$ and $\tilde{y}_{j}$ denote the corresponding hard labels and soft labels, respectively.
Cross-Optimizer Generalization.
In deep learning, models use various network architectures and optimization algorithms. Optimizers like SGD and Adam have unique properties that affect model performance and generalization. Therefore, evaluating a distilled dataset’s performance aross optimizers is essential to ensure its robustness across different training strategies. {definitionframe}
Definition 1 (Cross-optimizer Generalization) .
It refers to the capability of distilled datasets to maintain robust and consistent performance across different optimization algorithms.
@@ 3.2 Are Loss Functions Pulling the Strings in Synthetic Dataset Performance?
Why do we need to answer this question? Labels in dataset distillation are commonly and typically utilized through a variety of established loss functions, such as cross-entropy (CE) (Bridle, 1989), Kullback-Leibler (KL) divergence (Hinton et al., 2015), soft cross-entropy (Bridle, 1989) or mean squared error (MSE) (Nielsen, 2015). Specifically, the current state-of-the-art dataset distillation methods SRe2L (Yin et al., 2023) and RDED (Sun et al., 2024) employ KL divergence, DATM (Guo et al., 2024) uses soft cross-entropy, and G-VBSM (Shao et al., 2024) simultaneously utilizes MSE and CE. However, different distillation methods employ varying loss functions to train models on synthetic datasets, yet there is a notable lack of comprehensive comparison among these loss functions. Hence, we investigate the performance of four state-of-the-art dataset distillation methods under three commonly used loss functions.
Experiment settings. Our experiments span both small-scale Tiny-ImageNet and large-scale ImageNet-1K. Note that the synthetic dataset for DATM on ImageNet-1K is unavailable, preventing comparative analysis on this dataset. We conduct evaluations on these synthetic datasets with $\texttt{IPC}\in\{1,10,50\}$. Additional visualizations for Tiny-ImageNet and ImageNet-1K at $\texttt{IPC}=1$ and $\texttt{IPC}=50$ are provided in Appendix D. Notably, our empirical study is conducted from the end-user perspective: we treat the distillation of the synthetic dataset as a black box and apply different loss functions during the evaluation of the synthetic datasets.
Results and Analysis. The results, visualized in Figure 1, reveal that even for the same synthetic dataset, the evaluation model performance varies significantly under different loss functions. For example, the performance of SRe2L decreases by 13.2% on Tiny-ImageNet when switching from KL loss to soft CE loss but improves by 0.7% when using MSE+CE loss. These findings illustrate that the performance of models trained on synthetic datasets is highly sensitive to the choice of the loss function, highlighting the necessity for a unified and effective loss function in dataset distillation.
@@ 4 Method
The refined soft label is defined as $\tilde{y}_{j}\leftarrow\gamma\cdot\frac{y_{j}}{\|y_{j}\|}+(1-\gamma)\cdot\frac{\tilde{y}_{j}}{\|\tilde{y}_{j}\|}$, where $j$ is the j-th synthetic images, $y_{j}$ is the smoothed label obtained via label smoothing technique for hard label, and $\tilde{y}_{j}$ is the soft label. In our experiments, $\gamma=0.1$ is validated to be optimal through experiments depicted in Figure 2 in Section 5.6, and Figure 8 and Figure 9 in Appendix D.
@@
Prior study (Sun et al., 2024) points that representation learning from any samples $X$ to targets $Y$ is based on maximizing their mutual information. They propose to distill the dataset by maximizing $I_{\mathcal{V}}(X,Y)$, where $I_{\mathcal{V}}$ denotes the $\mathcal{V}$-information (Xu et al., 2020), which has demonstrated superior performance. However, we observe that most prior studies, including (Sun et al., 2024), have not investigated training models using $I_{\mathcal{V}}(X,Y)$. To address this gap, we explore the application of $I_{\mathcal{V}}(X,Y)$ by deriving an upper bound for the $\mathcal{V}$-information, as presented in Theorem 1 (Detailed proof is provided in Appendix A). {theoremframe}
Theorem 1 .
@@
where $\tau$ denotes the temperature parameter, $\mathcal{L}_{\textup{InfoNCE}}$ (Oord et al., 2018) serves as a proxy for $I_{\mathcal{V}}(X,Y)$ (Sun et al., 2024), and $f$ represents the similarity function.
Moreover, the targets $Y$ in the synthetic dataset are pre-generated using pre-trained models (Sun et al., 2024; Yin et al., 2023). These targets can be considered high-dimensional vectors that are approximately orthogonal to each other (Ma et al., 2022; Yu et al., 2023; Awasthi et al., 2024). Consequently, the term ${\mathbb{E}}[\boldsymbol{\phi}_{\boldsymbol{\theta}}(\mathbf{x}_{i})\cdot y_{j}]$ approaches zero, allowing (1) to be simplified as:
@@ 5 Experiments
In this section, we evaluate the superiority of our proposed GIFT across various datasets and architectures. First, we demonstrate the superior improvements of GIFT over the state-of-the-art dataset distillation in Section 5.2 and knowledge distillation methods in Section 5.3. Subsequently, we validate that GIFT achieves near-zero computational cost relative to baseline approaches ( Section 5.4). Additionally, we show that GIFT significantly improves cross-architecture and cross-optimizer generalization ( Section 5.5). To further understand the contributions of individual components, we perform detailed ablation studies ( Section 5.6). Finally, we demonstrate the effectiveness of GIFT in enhancing continual learning performance, as detailed in Appendix D. For additional experimental details and comprehensive results, refer to Appendix C and Appendix D.